OpenStudy (xmando98):

A binomial experiment consists of four independent trials. The probability of success in each trial is 9⁄25 . Find the probabilities of obtaining exactly 0 successes, 1 success, 2 successes, 3 successes, and 4 successes, respectively, in this experiment. a) [0.1678, 0.0944, 0.3716, 0.0299, 0.0168] b) [0.1678, 0.0105, 0.1593, 0.1194, 0.0168] c) [0.1678, 0.3775, 0.3185, 0.1194, 0.0168] d) [0, 0.1678, 0.3775, 0.3185, 0.1194] e) [0.1678, 0.0944, 0.3185, 0.3318, 0.0168] f) None of the above.

OpenStudy (xmando98):

please help me...

OpenStudy (eliesaab):

The probability of obtaining k successes in n trials with probability of success p in each trial is given by the formula \[ \binom{n}{k} p^k (1-p)^{n-k}=\frac{n!}{k!(n-k)!} p^k (1-p)^{n-k} \]

OpenStudy (eliesaab):

Here n=4, p=9/25, k=0,1,2,3,4 You have all what you need to compute 5 numbers, for k=0,1,2,3,4 Compute them and see which of your five results match one of the choices

OpenStudy (eliesaab):

For k=0; \[ \frac {4!}{0!(4-0)!}(9/25)^0(1-9/25)^4=(9/25)^0(16/25)^4=0.167772 \]

OpenStudy (eliesaab):

For k=1; you get \[ 4 (9/25)^1 (16/25)^3=0.377487 \]

OpenStudy (eliesaab):

Do k=3, 4, and 5 yourself and you will be done

OpenStudy (eliesaab):

Of course, you can use your calculator

OpenStudy (eliesaab):

I meant k=2,3 and 4 above